OpenStudy (hxkage):

For what value of p and q will the following pair of linear equations have infinitely many solutions? 4x + 5y = 2; (2p + 7q)x + (p+8q)y = 2p- q + 1

OpenStudy (hxkage):

@zepdrix

zepdrix (zepdrix):

They'll have infinitely many solutions `when they are both the same line`. That means in the second equation, we need `the stuff on x` to be 4, ya?

zepdrix (zepdrix):

\[\large\rm 2p+7q=4\]

zepdrix (zepdrix):

So then `the stuff on y`, needs to be how much?

OpenStudy (hxkage):

@zepdrix Sorry, I lost connection ;-;

zepdrix (zepdrix):

Oh I forgot I had OpenStudy open XD haha glad it gave me the orange notice

OpenStudy (hxkage):

5, the stuff on y should be 5

zepdrix (zepdrix):

\[\large\rm p+8q=5\]k good.

OpenStudy (hxkage):

xD

OpenStudy (hxkage):

Ty ty

zepdrix (zepdrix):

And the stuff on the right should be 2,\[\large\rm 2p- q + 1=2\]But we really only need two equations to solve this, so use whichever two are the easiest.

zepdrix (zepdrix):

\[\large\rm 2p+7q=4\]\[\large\rm p+8q=5\]Remember how to solve a system like this? :o

OpenStudy (hxkage):

Yeah, there are many methods ;-;

OpenStudy (hxkage):

Gimme a min, I'll solve it

OpenStudy (hxkage):

I got q = 6/9

zepdrix (zepdrix):

Same... hmm these are weird numbers :p

OpenStudy (hxkage):

Yup, quite weird ;-;

OpenStudy (hxkage):

P = 24/63

OpenStudy (hxkage):

Im not so sure about p

zepdrix (zepdrix):

Hmm thinking

zepdrix (zepdrix):

For infinite solutions, we need the same y-intercept and slope. Everything needs to be the same. So all three of these equations must hold true,\[\large\rm 2p+7q\qquad=4\]\[\large\rm ~~p+8q\qquad=5\]\[\large\rm 2p-~~q + 1=2\] Using only the first two equations, you should end up with \(\large\rm q=\frac23\) and \(\large\rm p=-\frac13\). But this doesn't hold true for the other equation :(( Hmmm. Is `no solutions` an option for your question?

zepdrix (zepdrix):

If you replace the p's with x's, and the q's with y's, then the system of coefficients looks like this,\[\large\rm 2x+7y\qquad=4\]\[\large\rm ~~x+8y\qquad=5\]\[\large\rm 2x-~~y + 1=2\]If you graph this system it looks like this, https://www.desmos.com/calculator/paysje4onr Notice that there is no `point` where all three lines intersect? Ya no solution :(

OpenStudy (hxkage):

Nope, they are asking for the values which will make it have infinite solutions

zepdrix (zepdrix):

There are no values. I don't think I made a mistake... but it's possible. Lemme see if someone smart is around >.< hmm

zepdrix (zepdrix):

Grr noone is online right now D:

OpenStudy (hxkage):

;-;

OpenStudy (hxkage):

It's cool, mate. I'll wait for somebody. Ty for your efforts! I appreciate it :) @zepdrix

OpenStudy (hxkage):

@Divyankasc A lil help, sis? ;-;

OpenStudy (hxkage):

I'm assuming we gotta do a1/a2 = b1/b2 and then cross multiply?

OpenStudy (divyankasc):

Yep

OpenStudy (hxkage):

So I did that and I got 6p= -3q

OpenStudy (hxkage):

so p = -3q/6

OpenStudy (hxkage):

Then substitute in the b1/b2 = c1/c2 part, I guess @Divyankasc

OpenStudy (divyankasc):

4x + 5y = 2; (2p + 7q)x + (p+8q)y = 2p- q + 1 For infinitely many solutions, a1/a2 = b1/b2 = c1/c2 a1/a2 = b1/b2 → 4/2p + 7q = 5/p + 8q 4(p + 8q) = 5(2p + 7q) 4p + 32q = 10p + 35q 6p = -3q q = -2p ....(1) a1/a2 = c1/c2 → 4/2p + 7q = 2/2p - q +1 4(2p - q + 1) = 2(2p + 7q) 4p - 4q + 1 = 4p + 14q 18q = 1 q = 1/18 From (1), 1/18 = -2p p =-1/36 Please check it again ↑↑↑ Because I do many mistakes while solving

OpenStudy (divyankasc):

Mistake! Wait

OpenStudy (hxkage):

Okiee

OpenStudy (divyankasc):

4x + 5y = 2; (2p + 7q)x + (p+8q)y = 2p- q + 1 For infinitely many solutions, a1/a2 = b1/b2 = c1/c2 a1/a2 = b1/b2 → 4/2p + 7q = 5/p + 8q 4(p + 8q) = 5(2p + 7q) 4p + 32q = 10p + 35q 6p = -3q q = -2p ....(1) a1/a2 = c1/c2 → 4/2p + 7q = 2/2p - q +1 4(2p - q + 1) = 2(2p + 7q) 4p - 4q + 4 = 4p + 14q 18q = 4 q = 4/18 = 2/9 From (1), 2/9= -2p p = -1/9

OpenStudy (divyankasc):

Does the answer match ?

OpenStudy (hxkage):

I don't have the answer key :/

OpenStudy (divyankasc):

Ohk, I think I amcorrect now

OpenStudy (divyankasc):

Then too, wait :P

OpenStudy (hxkage):

Gimme a min ;-;

OpenStudy (hxkage):

I'll check

OpenStudy (divyankasc):

Ahhh! I am really really sorry -_-

OpenStudy (divyankasc):

4x + 5y = 2; (2p + 7q)x + (p+8q)y = 2p- q + 1 For infinitely many solutions, a1/a2 = b1/b2 = c1/c2 a1/a2 = b1/b2 → 4/2p + 7q = 5/p + 8q 4(p + 8q) = 5(2p + 7q) 4p + 32q = 10p + 35q 6p = -3q q = -2p ....(1) a1/a2 = c1/c2 → 4/2p + 7q = 2/2p - q +1 4(2p - q + 1) = 2(2p + 7q) 8p - 4q + 4 = 4p + 14q 4p - 18q = -4 2p - 9q = -4 From (1), 2p + 18p = -4 20p = -4 p = -1/5 From (1), q = 2/5 This is the answer for sure!

OpenStudy (divyankasc):

I mean I am in so hurry that I did many mistakes

OpenStudy (hxkage):

Happens to me all the time! It's cool, Divya! Thank you for your help :)

OpenStudy (divyankasc):

And please make sure that you haven't written the question wrong. Recheck it.

OpenStudy (divyankasc):

NP :)

OpenStudy (hxkage):

The question's right, I checked it again. ;-;

OpenStudy (hxkage):

So is the answer, I believe. It looks good to me. c;

OpenStudy (divyankasc):

Cool, then my answer is also right.. Meow :3

OpenStudy (hxkage):

Yup.. :3

OpenStudy (hxkage):

You made a mistake. You wrote 4p - 18p = -4 then in the next line, u divided every term by 2 except -4.. @Divyankasc

OpenStudy (hxkage):

It's fine, Divya! :3

OpenStudy (divyankasc):

Ah! Okay ^_^" Good that you find out my mistake